OpenStudy (anonymous):

Write the expression 625^-1/4 in radical notation evaluate the radical expression

OpenStudy (zzr0ck3r):

o thats a negative sorry

OpenStudy (anonymous):

6(125) is not 625

OpenStudy (zzr0ck3r):

should have been a 5

OpenStudy (anonymous):

ok it should be a fraction not a root

OpenStudy (anonymous):

1/5

OpenStudy (zzr0ck3r):

im out of it

OpenStudy (zzr0ck3r):

sorry

OpenStudy (anonymous):

no worries thank you

OpenStudy (anonymous):

1/5 is equal to the correct answer but is not in the correct form

OpenStudy (zzr0ck3r):

\[1/\sqrt[4]{625} = 1/\sqrt[4]{5*5*5*5}= 1/5\]

OpenStudy (zzr0ck3r):

sure it is they wanted you to start with radical notation

OpenStudy (anonymous):

ok now it says evaluate the radical expression that would be the 1/5 correct

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

thank you so much for sticking around and helping me with it got a few more coming

OpenStudy (zzr0ck3r):

just make new posts. np